A random sample of 10 subjects have weights with a standard deviation of 11.8433 kg. What is the variance of their​ weights? Be sure to include the appropriate units with the result. The variance of the sample data is ? ▼ kg squared . kg. cubed . ​(Round to four decimal places as​ needed.)
Answer #1

From the given

Sample size =n =10

X=weight of a subject

Sample standard deviation

S=11.8433 kg

Varience of the weights

S2 = S*S = 11.8433*11.8433 = 140.2647 kg2​​​​​​

The varience of the sample data is 140.2647 kg squared.